# Importance sampling for Bayesian inference: polynomial-dimension dependent error bounds

### 02 Apr 2026 v1

@@ 4.3 Uniformly bounded observation function
The supremum of the observation magnitude, $M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{\mathbf{x}\in{\mathcal{X}}}\|h(\mathbf{x})\|$, characterizes the maximum signal energy the system captures as a function of the state dimension $d_{x}$. Under the assumptions of local saturation and uniform sensitivity, the forward map satisfies
$M(d_{x})\leq\sqrt{d_{y}}AM_{\sigma}d_{x}.$ (45)
@@
$|h_{i}(\mathbf{x})|\leq\sum_{j=1}^{d_{x}}|a_{ij}||\sigma(x_{j})|\leq\sum_{j=1}^{d_{x}}AM_{\sigma}=AM_{\sigma}d_{x}.$ (46)
Taking the Euclidean norm of the vector $h(\mathbf{x})=(h_{1}(\mathbf{x}),\dots,h_{d_{y}}(\mathbf{x}))^{\top}$, we obtain
$\|h(\mathbf{x})\|=\left(\sum_{i=1}^{d_{y}}|h_{i}(\mathbf{x})|^{2}\right)^{1/2}\leq\left(\sum_{i=1}^{d_{y}}(AM_{\sigma}d_{x})^{2}\right)^{1/2}=\sqrt{d_{y}}AM_{\sigma}d_{x}.$ (47)
Since this bound holds for all $\mathbf{x}\in{\mathcal{X}}$, the result follows. This formulation ensures that for any finite dimension $d_{x}$, the observation operator remains bounded on the domain $\mathcal{X}$}. Notably, the growth of the observation map satisfies $M(d_{x})\leq P(d_{x})\mathrel{\mathop{\ordinarycolon}}=a_{1}d_{x}$ for $a_{1}=\sqrt{d_{y}}AM_{\sigma}>0$, where $P$ is a polynomial of degree 1 in the state dimension. Consequently, for any observation model $Y=h(X)+V$ where the noise $V$ admits a polynomial radial profile as described in Remark 8 $M(d_{x})\propto d_{x}$ grows linearly and satisfies the conditions of Corollary 4.2 (item 2). This ensures that the requirements of Theorem 3.3 are fulfilled, guaranteeing that the error of the IS estimator $\pi_{Y}^{N}$ scales at most polynomially (in fact, linearly) with the state dimension $d_{x}$.
@@ 5.2 Importance sampling for Bayesian inference
To achieve $\mathrm{MSE}\leq\varepsilon$, one needs $N\gtrsim\varepsilon^{-1}\exp\{\tau\}$. Thus the computational cost can grows exponentially with the intrinsic dimension.
@@
However, instead of estimating $\rho_{Y}$ pointwise in $Y$, we integrate with respect to the law of $Y$ and exploit structural properties of the model. For the linear–Gaussian setting, Corollary 3.4 combined with (32) yields the uniform bound
$\|\pi_{Y}(f)-\pi_{Y}^{N}(f)\|_{2}\leq\frac{B_{p}\|f\|_{\infty}}{\left((2^{2/3}\pi)^{d_{y}}|R|\right)^{3/2}}\frac{1}{\sqrt{N}}.$
All constants are independent of the state dimension $d_{x}$, and the model dependence enters only through the observation noise covariance $R$ and the observation dimension $d_{y}$. In particular, if the eigenvalues of $R$ are sufficiently large (see Section 4.1.1), the bound remains finite, and may even decrease, as $d_{y}\to\infty$.
@@ 6 Conclusions
We have investigated the approximation errors of a random importance sampling algorithm for numerical Bayesian inference in high-dimensional settings. The analysis is carried out in a probabilistic framework in which the observations are treated as random variables, and the posterior distribution is therefore itself a random measure. This setting allows us to exploit the probabilistic structure of the model and to formulate error bounds in expectation, rather than conditioning on a fixed realization of the data, and analyse the specifically, worst-case scenario.
Let $\pi_{Y}$ and $\pi_{Y}^{N}$ denote the true posterior distribution of a $d_{x}$-dimensional r.v. $X$ given an observation $Y$ and its IS approximation, respectively. In full generality, classical upper bounds for the $L^{2}$ approximation error $|\pi_{y}(f)-\pi_{y}^{N}(f)|$, for a given realization $Y=y$ and for any bounded test functions $f$, may grow exponentially with dimension, in agreement with existing results for IS and related Monte Carlo schemes [4, 35, 36, 1, 9].
In this paper, we have introduced a refined analysis, treating observations as r.v.s, we relate the approximation error $|\pi_{Y}(f)-\pi_{Y}^{N}(f)|$ to the Bochner integrability of a suitable link function $\ell^{\,d_{x}}$, which generates normalized likelihoods for the model. We have shown that the $L^{2}$ error bounds are finite and converge at the standard Monte Carlo rate ${\mathcal{O}}(N^{-1/2})$, for arbitrarily large dimension, if and only if the link function is Bochner integrable. This result provides a precise and explicit characterization of when high dimensionality does not lead to a breakdown of the algorithm.
Moreover, the dependence of the error bounds on dimension can be quantified in terms of the moments of the link function with respect to the marginal distribution of the observations. This perspective enables the identification of sufficient conditions under which the $L^{1}$ and $L^{2}$ error bounds grow only polynomially with dimension, as well as conditions under which they remain uniformly bounded. These results provide a rigorous explanation of how the curse of dimensionality can be alleviated for some families of models.
We have illustrated the general theory through several concrete examples. In particular, for linear and Gaussian models, we have proved analytically that Bochner $2$-integrability of the link function alone is sufficient to ensure uniform $L^{2}$ error bounds, independently of dimension. We have also considered models with bounded or dominated observation functions, conditions that are commonly satisfied in practical applications, for instance when sensors operate within a finite dynamic range. In all cases, we provide explicit analytical arguments showing how the structure of the model leads to favorable dimensional scaling of the average errors.
@@ Appendix A Use of importance functions
$\widetilde{g}_{y}^{\,d_{x}}(x)\mathrel{\mathop{\ordinarycolon}}=g_{y}^{\,d_{x}}(x)\,\rho(x).$ (54)
@@
Using the definitions in (53)–(55), it is straightforward to verify that the procedure described above coincides exactly with Algorithm 1. Consequently, to assess the performance of Algorithm 1 for the model $\mathcal{M}$}, it suffices to analyze the performance of Algorithm 2 for the modified model $\widetilde{{\mathcal{M}}}$, that is, to study the Bochner integrability of the associated link function $\widetilde{\ell^{\,d_{x}}}$ under $\widetilde{{\mathcal{M}}}$. To this end, it is also necessary to ensure that the likelihood function $\widetilde{g}_{y}^{\,d_{x}}$ of the modified model $\widetilde{{\mathcal{M}}}$ satisfies $\widetilde{g}_{y}^{\,d_{x}}\in L^{2}(\nu)$ fora all $y\in{\mathcal{Y}}$.
@@ Appendix B Statistical analysis of the normalization constant
$Z\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{X}}g_{y}^{d_{x}}(x)\,\pi_{0}({\sf d}x)=\pi_{0}(g_{y}^{d_{x}}).$ (56)
@@
$\mathrm{Var}(\hat{Z}_{N})=\frac{1}{N}\left(\mathbb{E}_{\pi_{0}}\!\left[(g_{y}^{d_{x}}(X))^{2}\right]-(\mathbb{E}_{\pi_{0}}[g_{y}^{d_{x}}(X)])^{2}\right).$ (58)
@@
$\mathrm{Var}(\hat{Z}_{N})=\frac{1}{N}\left(\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{2}-Z^{2}\right).$ (59)
@@
$\frac{\mathrm{Var}(\hat{Z}_{N})}{Z^{2}}=\frac{1}{N}\left(\frac{\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{2}}{Z^{2}}-1\right).$ (60)
@@ C.1 Preliminaries
$\left|\pi_{y}(f)-\pi_{y}^{N}(f)\right|\leq\frac{1}{\mathtt{c}\,\pi_{0}(g_{y}^{d_{x}})}\left[\left|\pi_{0}^{N}(fl_{y}^{d_{x}})-\pi_{0}(fl_{y}^{d_{x}})\right|+\|f\|_{\infty}\left|\pi_{0}(l_{y}^{d_{x}})-\pi_{0}^{N}(l_{y}^{d_{x}})\right|\right].$ (62)
@@ D.2 Boundedness of expectation functionals
$=K_{2}^{d_{z}}\|\upsilon\|_{L^{\infty}({\mathcal{Y}};L^{2}(\pi_{0}))}\|\varsigma\|_{L^{\infty}({\mathcal{Y}};L^{2}(\pi_{0}))}<\infty.$
@@
$\left|\pi_{Y}(f)-\pi_{Y}^{N}(f)\right|\leq\frac{1}{\pi_{0}(l_{Y}^{d_{x}})}\left[\left|\pi_{0}(f\,l_{Y}^{d_{x}})-\pi_{0}^{N}(f\,l_{Y}^{d_{x}})\right|+\|f\|_{\infty}\left|\pi_{0}^{N}(l_{Y}^{d_{x}})-\pi_{0}(l_{Y}^{d_{x}})\right|\right].$ (66)
@@
Appendix G Elliptically symmetric likelihood families
G.1 Mahalanobis-distance geometry

### 28 May 2026 v2

@@ 4.3.1 Example: Polynomial growth of $M(d_{x})$
The supremum of the observation magnitude, $M(d_{x})\mathrel{\mathop{\ordinarycolon}}=\sup_{\mathbf{x}\in{\mathcal{X}}}\|h^{d_{x}}(\mathbf{x})\|$, characterizes the maximum signal energy the system captures as a function of the state dimension $d_{x}$. Under the assumptions of local saturation and uniform sensitivity, the forward map satisfies
$M(d_{x})\leq\sqrt{d_{y}}AM_{\sigma}d_{x}.$
@@
$|h_{i}^{d_{x}}(\mathbf{x})|\leq\sum_{j=1}^{d_{x}}|a_{ij}||\sigma(x_{j})|\leq\sum_{j=1}^{d_{x}}AM_{\sigma}=AM_{\sigma}d_{x}.$
Taking the Euclidean norm of the vector $h^{d_{x}}(\mathbf{x})=(h_{1}^{d_{x}}(\mathbf{x}),\dots,h_{d_{y}}^{d_{x}}(\mathbf{x}))^{\top}$, we obtain
$\|h^{d_{x}}(\mathbf{x})\|=\left(\sum_{i=1}^{d_{y}}|h_{i}^{d_{x}}(\mathbf{x})|^{2}\right)^{1/2}\leq\left(\sum_{i=1}^{d_{y}}(AM_{\sigma}d_{x})^{2}\right)^{1/2}=\sqrt{d_{y}}AM_{\sigma}d_{x}.$
Since this bound holds for all $\mathbf{x}\in{\mathcal{X}}$, the result follows. This formulation ensures that for any finite dimension $d_{x}$, the observation operator remains bounded on the domain $\mathcal{X}$}. Notably, the growth of the observation map satisfies $M(d_{x})\leq P(d_{x})\mathrel{\mathop{\ordinarycolon}}=a_{1}d_{x}$ for $a_{1}=\sqrt{d_{y}}AM_{\sigma}>0$, then $P$ is a polynomial of degree 1 in the state dimension. Consequently, for any observation model $Y=h^{d_{x}}(X)+V$ where the noise $V$ admits a polynomial radial profile as described in Remark 9 $M(d_{x})\propto d_{x}$ grows linearly and satisfies the conditions of Corollary 4.2 (item 2). This ensures that the requirements of Theorem 3.3 are fulfilled, guaranteeing that the error of the IS estimator $\pi_{Y}^{N}$ scales at most polynomially (in fact, linearly) with the state dimension $d_{x}$.
@@ 5.2 Importance sampling for Bayesian inference
To achieve $\mathrm{MSE}\leq\varepsilon$, one needs $N\gtrsim\varepsilon^{-1}\exp\{\tau\}$. Thus the computational cost can grow exponentially with the intrinsic dimension.
@@
However, instead of estimating $\rho_{Y}$ pointwise in $Y$, we integrate with respect to the law of $Y$ and exploit structural properties of the model. In the linear–Gaussian setting, assuming the observation dimension $d_{y}$ is fixed, Corollary 3.4, together with (38) and (39), yields the uniform bound
$\|\pi_{Y}(f)-\pi_{Y}^{N}(f)\|_{2}\leq\left(1+\frac{C_{\Sigma}\,C_{A}^{2}}{\lambda_{d_{y}}(R)}\right)^{d_{y}}\frac{B_{p}\|f\|_{\infty}}{\sqrt{N}}.$
All constants are independent of the state dimension $d_{x}$, and the model dependence enters only through the observation noise covariance $R$ and the observation dimension $d_{y}$.
@@ 6 Conclusions
We have investigated the approximation errors of importance samplers for Bayesian inference in high-dimensional settings. The analysis is carried out in a probabilistic framework in which the observations are treated as r.v.s, and the posterior distribution is itself a random measure. This setting allows us to exploit the probabilistic structure of the model and to formulate expected error bounds, rather than conditioning on a fixed realization of the data and analyse the worst-case scenario.
Let $\pi_{Y}$ and $\pi_{Y}^{N}$ denote the true posterior distribution of a $d_{x}$-dimensional r.v. $X$ given an observation $Y$ and its IS approximation, respectively. In full generality, classical upper bounds for the $L^{2}$ approximation error $\|\pi_{y}(f)-\pi_{y}^{N}(f)\|_{2}$, for a given realization $Y=y$ and for any bounded test functions $f$, may grow exponentially with dimension, in agreement with existing results for IS and related Monte Carlo schemes [5, 36, 37, 1, 10].
In this paper, we have introduced a refined analysis, treating observations as r.v.s, we relate the approximation error $\|\pi_{Y}(f)-\pi_{Y}^{N}(f)\|_{2}$ to the Bochner integrability of a suitable link function $\ell^{\,d_{x}}$, which generates normalized likelihoods for the model. We have shown that the $L^{2}$ error bounds are finite and converge at the standard Monte Carlo rate ${\mathcal{O}}(N^{-1/2})$, for arbitrarily large dimension, if and only if the link function is Bochner integrable. This result provides a precise and explicit characterization of when high dimensionality does not lead to a breakdown of the algorithm.
Moreover, the dependence of the error bounds on the dimension $d_{x}$ can be quantified in terms of the moments of the link function w.r.t. the marginal distribution of the observations. This perspective enables the identification of sufficient conditions under which the $L^{1}$ and $L^{2}$ error bounds grow only polynomially with dimension, as well as conditions under which they remain uniformly bounded. These results provide a rigorous explanation of how some families of models do not suffer from a curse of dimensionality.
We have illustrated the general theory through several examples. In the linear Gaussian setting, we make explicit how the Bochner $2$-integrability constant depends on the model parameters, and how its growth w.r.t. the state dimension $d_{x}$ is polynomial under regularity conditions. In particular, by appropriately bounding the singular values of the observation operator and the eigenvalues of the state covariance matrix, we show that it is possible to obtain error bounds that remain uniform regardless of the state dimension $d_{x}$. We have also considered models with bounded or dominated observation function, conditions that are commonly satisfied in practical applications, for instance when physical sensors operate within a finite dynamic range. In all cases, we provide explicit analytical arguments showing how the structure of the model leads to favorable dimensional scaling of the average errors w.r.t. the observation distribution.
@@ Appendix A Use of importance functions
$\widetilde{g}_{y}^{\,d_{x}}(x)\mathrel{\mathop{\ordinarycolon}}=g_{y}^{\,d_{x}}(x)\,\rho(x).$
@@
Using the definitions in (50)–(51), it is straightforward to verify that the procedure described above coincides exactly with Algorithm 1. Consequently, to assess the performance of Algorithm 1 for the model $\mathcal{M}$}, it suffices to analyze the performance of Algorithm 2 for the modified model $\widetilde{{\mathcal{M}}}$, that is, to study the Bochner integrability of the associated link function $\widetilde{\ell^{\,d_{x}}}$ under $\widetilde{{\mathcal{M}}}$. To this end, it is also necessary to ensure that the likelihood function $\widetilde{g}_{y}^{\,d_{x}}$ of the modified model $\widetilde{{\mathcal{M}}}$ satisfies $\widetilde{g}_{y}^{\,d_{x}}\in L^{2}(\nu)$ for all $y\in{\mathcal{Y}}$.
@@ Appendix B Statistical analysis of the normalization constant
$Z\mathrel{\mathop{\ordinarycolon}}=\int_{\mathcal{X}}g_{y}^{d_{x}}(x)\,\pi_{0}({\sf d}x)=\pi_{0}(g_{y}^{d_{x}}).$
@@
$\mathrm{Var}(\hat{Z}_{N})=\frac{1}{N}\left(\mathbb{E}_{\pi_{0}}\!\left[(g_{y}^{d_{x}}(X))^{2}\right]-(\mathbb{E}_{\pi_{0}}[g_{y}^{d_{x}}(X)])^{2}\right).$
@@
$\mathrm{Var}(\hat{Z}_{N})=\frac{1}{N}\left(\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{2}-Z^{2}\right).$
@@
$\frac{\mathrm{Var}(\hat{Z}_{N})}{Z^{2}}=\frac{1}{N}\left(\frac{\|g_{y}^{d_{x}}\|_{L^{2}(\pi_{0})}^{2}}{Z^{2}}-1\right).$
@@ C.1 Preliminaries
$\left|\pi_{y}(f)-\pi_{y}^{N}(f)\right|\leq\frac{1}{\mathtt{c}\,\pi_{0}(g_{y}^{d_{x}})}\left[\left|\pi_{0}^{N}(fl_{y}^{d_{x}})-\pi_{0}(fl_{y}^{d_{x}})\right|+\|f\|_{\infty}\left|\pi_{0}(l_{y}^{d_{x}})-\pi_{0}^{N}(l_{y}^{d_{x}})\right|\right].$
@@ D.2 Boundedness of expectation functionals
$=K_{2}^{d_{x}}\|\upsilon\|_{L^{\infty}({\mathcal{Y}};L^{2}(\pi_{0}))}\|\varsigma\|_{L^{\infty}({\mathcal{Y}};L^{2}(\pi_{0}))}<\infty.$
@@
$\left|\pi_{Y}(f)-\pi_{Y}^{N}(f)\right|\leq\frac{1}{\pi_{0}(l_{Y}^{d_{x}})}\left[\left|\pi_{0}(f\,l_{Y}^{d_{x}})-\pi_{0}^{N}(f\,l_{Y}^{d_{x}})\right|+\|f\|_{\infty}\left|\pi_{0}^{N}(l_{Y}^{d_{x}})-\pi_{0}(l_{Y}^{d_{x}})\right|\right].$
@@
Appendix G Numerical example
To illustrate the dimensional behavior predicted by the theoretical bounds developed throughout the paper, we consider a family of linear Gaussian models for which the relevant spectral quantities can be controlled explicitly as the state dimension $d_{x}$ varies.
We present two numerical examples, both with fixed observation dimension $d_{y}$, but with different spectral scalings. In the first example, the model is normalized so that the quantity
$\sigma_{1}^{2}(A^{d_{x}})\lambda_{1}(\Sigma_{x}^{d_{x}})$
grows polynomially with respect to $d_{x}$. As predicted by the theory, this produces polynomial growth of the Bochner constant $K_{2}^{d_{x}}$ and of the corresponding importance sampling error.
In the second example, the normalization is chosen so that the previous spectral quantity remains uniformly bounded with respect to $d_{x}$. In this regime, the theoretical bounds predict that both the Bochner constant and the importance sampling error remain uniformly stable as the state dimension increases.
Together, these examples illustrate how the spectral structure of the model determines the dimensional scaling of the average errors with respect to the observation distribution.
Throughout both examples, we fix
$d_{y}=3,\qquad R=I_{d_{y}},$
and, for each
$d_{x}\in\{5,10,20,50,10^{2},5\cdot 10^{2},10^{3},10^{4}\},$
we generate random matrices we generate random matrices
$\widetilde{A}^{d_{x}}\in\mathbb{R}^{d_{y}\times d_{x}},\qquad\widetilde{B}^{d_{x}}\in\mathbb{R}^{d_{x}\times d_{x}},$
whose entries are independent and uniformly distributed in $(0,1)$. Since $\sigma_{1}(\widetilde{A}^{d_{x}})=\|\widetilde{A}^{d_{x}}\|_{2}$, this assumption is enough to ensure that $\sigma_{1}(\widetilde{A}^{d_{x}})>0$. We then define
$\widetilde{\Sigma}_{x}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\widetilde{B}^{d_{x}}(\widetilde{B}^{d_{x}})^{\top}+\eta I_{d_{x}},~~\text{with}~~\eta=10^{-2}.$
Thus, $\widetilde{\Sigma}_{x}^{d_{x}}$ is symmetric positive definite. Indeed, for every $x\neq 0$,
$x^{\top}\widetilde{\Sigma}_{x}^{d_{x}}x=x^{\top}\widetilde{B}^{d_{x}}(\widetilde{B}^{d_{x}})^{\top}x+\eta x^{\top}x=\|(\widetilde{B}^{d_{x}})^{\top}x\|^{2}+\eta\|x\|^{2}.$
Since $\eta\|x\|^{2}>0,$ we obtain
$x^{\top}\widetilde{\Sigma}_{x}^{d_{x}}x>0,\qquad\forall x\neq 0.$
Hence, $\widetilde{\Sigma}_{x}^{d_{x}}\succ 0.$
G.1 Polynomial scaling
The model matrices are obtained through the spectral normalizations
$A^{d_{x}}=d_{x}^{1/4}\frac{\widetilde{A}^{d_{x}}}{\sigma_{1}(\widetilde{A}^{d_{x}})},\qquad\Sigma_{x}^{d_{x}}=d_{x}^{1/2}\frac{\widetilde{\Sigma}_{x}^{d_{x}}}{\lambda_{1}(\widetilde{\Sigma}_{x}^{d_{x}})}.$ (58)
By (58) and the homogeneity of singular values and eigenvalues, namely
$\sigma_{1}(cM)=c\,\sigma_{1}(M),\qquad\lambda_{1}(cM)=c\,\lambda_{1}(M),\qquad c>0,$
we obtain the exact identities
$\sigma_{1}(A^{d_{x}})=d_{x}^{1/4},\qquad\lambda_{1}(\Sigma_{x}^{d_{x}})=d_{x}^{1/2}.$
Consequently,
$\sigma_{1}^{2}(A^{d_{x}})\lambda_{1}(\Sigma_{x}^{d_{x}})=d_{x}.$
This quantity is the key mechanism behind the polynomial behavior observed below. Indeed, applying the upper bound in (37) to
$K_{2}^{d_{x}}\leq\left(1+\frac{\lambda_{1}(\Sigma_{x}^{d_{x}})}{\lambda_{d_{y}}(R)}\sigma_{1}^{2}(A^{d_{x}})\right)^{d_{y}}$
yields
$K_{2}^{d_{x}}\leq(1+d_{x})^{d_{y}}=\mathrel{\mathop{\ordinarycolon}}P(d_{x}).$ (59)
For each value of $d_{x}$, we choose a prior mean
$\mu_{x}^{d_{x}}\in\mathbb{R}^{d_{x}},$
which is generated for the simulation as a sample from a standard $d_{x}$-dimensional Gaussian distribution. We then sample a state
$x^{\star}\sim\mathcal{N}(\mu_{x}^{d_{x}},\Sigma_{x}^{d_{x}}),$
and generate one observation
$y=A^{d_{x}}x^{\star}+v,\qquad v\sim\mathcal{N}(0,I_{d_{y}}).$
Since the model is linear and Gaussian, the posterior distribution of $X$ is available in closed form. In particular, the posterior mean is
$m_{y}^{d_{x}}=\mu_{x}^{d_{x}}+\Sigma_{x}^{d_{x}}[A^{d_{x}}]^{\top}\left(A^{d_{x}}\Sigma_{x}^{d_{x}}[A^{d_{x}}]^{\top}+I_{d_{y}}\right)^{-1}\left(y-A^{d_{x}}\mu_{x}^{d_{x}}\right).$
We compare this exact posterior mean with an importance sampling approximation with $N=400$ particles sampled from the prior,
$x^{1},\ldots,x^{N}\sim\mathcal{N}(\mu_{x}^{d_{x}},\Sigma_{x}^{d_{x}}).$
For the realized observation $y$, the normalized weights are
$w_{i}(y)=\frac{\exp\left\{-\frac{1}{2}\|y-A^{d_{x}}x^{i}\|^{2}\right\}}{\sum_{j=1}^{N}\exp\left\{-\frac{1}{2}\|y-A^{d_{x}}x^{j}\|^{2}\right\}},\qquad i=1,\ldots,N,$
and the corresponding importance sampling approximation of the posterior mean is
$m_{y,N}^{d_{x}}=\sum_{i=1}^{N}w_{i}(y)x^{i}.$
For each dimension $d_{x}$, we compute the Euclidean error
${\mathcal{E}}_{N}^{d_{x}}\mathrel{\mathop{\ordinarycolon}}=\|m_{y,N}^{d_{x}}-m_{y}^{d_{x}}\|_{2},$
as well as the Bochner constant, which by Eq. (33) takes the form
$K_{2}^{d_{x}}=\frac{\mathinner{\!\left\lvert A^{d_{x}}\Sigma_{x}^{d_{x}}[A^{d_{x}}]^{\top}+I_{d_{y}}\right\rvert}}{\mathinner{\!\left\lvert I_{d_{y}}\right\rvert}}=\mathinner{\!\left\lvert A^{d_{x}}\Sigma_{x}^{d_{x}}[A^{d_{x}}]^{\top}+I_{d_{y}}\right\rvert}$.
Recall that, by (59)
$K_{2}^{d_{x}}\leq(1+d_{x})^{d_{y}},$
and therefore the theoretical upper bound grows almost polynomially. Recall that, in this example, the observation dimension is fixed and equal to $d_{y}=3$.
Figure 1 compares $K_{2}^{d_{x}}$, the importance sampling error ${\mathcal{E}}_{N}^{d_{x}}$, the reference polynomial $P(d_{x})=(1+d_{x})^{3}$, and the exponential reference $e^{d_{x}}$. The exponential curve is displayed only for $d_{x}\leq 50$, since it grows much faster than the polynomial quantities.
Figure 1: The Bochner constant $K_{2}^{d_{x}}$ and the importance sampling error ${\mathcal{E}}_{N}^{d_{x}}=\|m_{y,N}^{d_{x}}-m_{y}^{d_{x}}\|_{2}$ as a function of the state dimension $d_{x}$, for $d_{y}=3$ and $N=400$. The polynomial reference predicted $P(d_{x})=(1+d_{x})^{d_{y}}$ and the exponential reference $e^{d_{x}}$ are shown for comparison. The results illustrate the polynomial growth predicted by the theoretical bounds.
G.2 Uniform bound
If instead of allowing the spectral quantities to grow with $d_{x}$, we impose uniform bounds of the form
$\sigma_{1}(A^{d_{x}})\leq C_{A},\qquad\lambda_{1}(\Sigma_{x}^{d_{x}})\leq C_{\Sigma},\qquad\text{for all }d_{x},$
for some constants $C_{A},C_{\Sigma}>0$ independent of $d_{x}$, which can be achieved by a straightforward modification of the normalization in (58), then the key quantity governing the growth satisfies
$\sigma_{1}^{2}(A^{d_{x}})\lambda_{1}(\Sigma_{x}^{d_{x}})\leq C_{A}^{2}C_{\Sigma}.$
Consequently, the determinant bound in (37) yields
$K_{2}^{d_{x}}\leq\left(C_{A}^{2}C_{\Sigma}+1\right)^{d_{y}}\mathrel{\mathop{\ordinarycolon}}={\mathcal{K}}_{2},$
therefore, the Bochner constant remains uniformly bounded w.r.t. $d_{x}$ and the corresponding error of the importance sampling approximation, ${\mathcal{E}}_{N}^{d_{x}}$, remains uniformly bounded as well. This behavior is illustrated in Figure 2, where both quantities remain stable as $d_{x}$ increases, in contrast with the polynomial growth observed in Figure 1.
Figure 2: The Bochner constant $K_{2}^{d_{x}}$ and the importance sampling error ${\mathcal{E}}_{N}^{d_{x}}=\|m_{y,N}^{d_{x}}-m_{y}^{d_{x}}\|_{2}$ as a function of the state dimension $d_{x}$, for $d_{y}=3$ and $N=400$. The uniform theoretical bound $\mathcal{K}_{2}=(1+3)^{d_{y}}$ and the exponential reference $e^{d_{x}}$ are shown for comparison. The results illustrate the uniform behavior predicted by the theoretical bounds, where both the Bochner constant and the importance sampling error remain stable as the state dimension increases.
Appendix H Elliptically symmetric likelihood families
H.1 Mahalanobis-distance geometry